HUBBLE SPACE TELESCOPE DAILY REPORT #3103 PERIOD COVERED: 0000Z (UTC) 04/25/02 - 0000Z (UTC) 04/26/02 Daily Status Report as of 116/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 9058 (Trigonometric Parallax of the Globular Cluster 47 Tucanae) The WF/PC-2 was used measure the trigonometric parallax of 47 Tuc, which is about 1/4 milliarcsec, to an accuracy of 10%. By repeating a field that was observed in 1995, we will measure the secular motion of the cluster stars with respect to the SMC background with sufficient accuracy that the parallax can be measured directly from the displacements within Cycle 10. Constraints of orientation, and the shape of the WFPC2 footprint, require that we take images at 3-month intervals, rotating 90 degrees each time. The proposal completed with no reported problems. 1.2 Completed Six Sets of STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.3 Completed Two Sets of ACS/WFC 9575 (Default {Archival} Pure Parallel Program) The Advanced Camera for Surveys (WFC) was used to test ACS pure parallels in POMS. There were no reported problems. 1.4 Completed Ten Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used 1) to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.5 Completed Two Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.6 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.7 Completed Seven Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.8 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.9 Completed WF/PC-2 9050 (Outflow Collimation in Bipolar Symbiotic Nebulae) The WF/PC-2 was used to observe flow collimation in evolved stars that is neither expected nor understood. Classical theories of stellar evolution do not predict and cannot explain this bipolarity. More exotic concepts {binary interactions, spun-up atmospheres, poloidal or toroidal magnetized winds} have been proposed, but observations are yet to verify or falsify any of their predictions. This proposal will probe the near-nuclear morphology and kinematics of four bright, low-extinction targets whose large-scale structure is highly bipolar. The goal is to provide a detailed description of the circumnuclear outflows, to uncover the physical structure and nature of the collimator, and to evaluate the speculative collimation mechanisms. The bright nucleus has hampered efforts to explore the nebular collimators that lie close to the star, so we'll use STIS to disperse the nuclear light and, thus, to avoid its glare. A secondary goal is to obtain second-epoch WFPC2 images of all targets. There were no reported problems. 1.10 Completed Three Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.11 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) The Space Telescope Imaging Spectrograph (MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise in order to monitor the effects of thermal changes on the NUV dark rate. No problems occurred. 1.12 Completed STIS/CCD 9157 (Fundamental Properties Of L-Type Dwarfs In Binaries) The Space Telescope Imaging Spectrograph (CCD) was used to characterize the physical properties of eight L-dwarfs in four binary systems. The goal is to obtain astrometric, photometric and spectroscopic measurements of each component that will yield basic information on atmospheric and dynamical properties. The proposal completed nominally. 1.13 Completed Two Sets of WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. The proposal had no problems. 1.14 Completed ACS/WFC 9584 (ACS Default {Archival} Pure Parallel Program II) The Advanced Camera for Surveys (WFC) was used to test ACS pure parallels. The proposal completed with no reported problems. 1.15 Completed WF/PC-2 8940 (Cycle 10 Earth Flats) The WF/PC-2 was used to monitor flatfield stability by obtaining sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed without incident. 1.16 Completed Four Sets of NICMOS/1/2/3 8944 (Filter Wheel/Mechanisms Mini-Functional Test) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used to perform an early engineering test to verify the aliveness, functionality, operability, and electro-mechanical calibration of the NICMOS filter wheel motors and assembly. This was the first use of the NICMOS filter wheel mechanisms since they were disabled by ground command in January, 1999. This test was designed to obviate concerns over possible deformation or breakage of the fitter wheel "soda-straw" shafts due to excess rotational drag torque and/or bending moments which may be imparted due to changes in the dewar metrology from warm-up/cool-down. No anomalous incident occurred. 1.17 Completed STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) The Space Telescope Imaging Spectrograph (CCD) was used to perform a snapshot survey of all known LMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, and to probe the chemical enrichment history of the LMC. There were no reported problems. 1.18 Completed STIS/CCD/MA1 8962 (Contamination Monitor) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to observe a flux standard GRW +70 5824 to assess whether FUV sensitivity is significantly below performance obtained during execution of 8961 {End of BEA Test} and prior to SM3B. Significantly degraded sensitivity will trigger additional observations of flux standards to validate the initial result and/or track subsequent {de}contamination. GRW +70 5824 is observed bimonthly in the same mode as part of the regular STIS monitoring program. Measurement accuracy will be limited global sensitivity fluctuations {less than 5%}, rather than photon counting statistics. No anomalous activity was observed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 26 Successful: 25 Per HSTAR 8627, the roll delay update at 115/220401Z failed due to tracker #1. The subsequent full maneuver updates were successful. 2.3 Operations Notes: The NSSC-1 status buffer was dumped and reset a5 115/1116Z per ROP NS-3. Operations requests were used to: Enable Filter Move, Day 115-1 @115/1618z Enable Filter Wheel Move, Day 115-2 @115/2236z Vehicle Control Law Rate Gain Factor Change @115/1526z HLGBU GYRO Triplets U/L @115/1447z Configure SDH Science Tasking Orders @115/1610z Set Up Tasking Order for FSW Test @115/1616z Configure Tasking Order after FSW Test Activities @115/2304z A TTR was generated when there was a required re-transmits (at 115/190618Z and at 115/191520Z) during a NSSC-1 uplink. Accordingly, SI C&DH error were reset at 115/1917Z, using ROP NS-5. The 486 engineering status buffer limits were adjusted at 115/2213Z as directed by ROP DF-18A. Using ROP DF-01A, a 486 table was dumped at 115/1923Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW